OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet. Show your work The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6.

OpenStudy (anonymous):

each braclets $6

OpenStudy (anonymous):

the total net cost for 8 bracelets $48, where each bracelets $6

OpenStudy (anonymous):

1

OpenStudy (anonymous):

it is 1

OpenStudy (anonymous):

? right

OpenStudy (anonymous):

why 1?

OpenStudy (anonymous):

8x + 6 = 54

OpenStudy (anonymous):

olve for x to give cost of 1 bracelet

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

how would u set it up

OpenStudy (anonymous):

8x + 6 = 54 8x = 54 -6 8x = 48 x = 48/8 x = 6

OpenStudy (anonymous):

this enough???